# Sparse factorization of the square all-ones matrix of arbitrary order

### v1

@@ 6.2 Decentralized optimization
$\text{minimize}\ \ f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),$ (30)
@@
where the optimization variable is $x\in\mathbb{R}^{d}$, and each component function $f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable and potentially nonconvex. Each agent $i\in{\mathcal{V}}$ only has access to one component function $f_{i}$, and agents communicate with each other via (time-varying) topologies $\{{\mathcal{G}}^{(k)}\}$. It can be shown that the decentralized average consensus problem is a special case of the optimization Problem (30) with $f_{i}(x)=\tfrac{1}{2}\|x-x_{i}^{(0)}\|_{2}^{2}$.
@@
In the context of decentralized optimization, a sparse factorization of $J$ offers sequences of graphs that satisfy the finite-time consensus property, and incorporating such graph sequences in decentralized optimization algorithms could significantly reduce the per-iteration communication cost in the algorithm while achieving a comparable convergence rate (compared with decentralized algorithms using traditional communication protocols) [13, 7]. Also note that most existing analyses for decentralized optimization algorithms need the assumption that the weight (mixing) matrices $\{W^{(k)}\}$ are doubly stochastic, which is satisfied by the DSHB factor and the $T$-factor (21) in the SDS factorization. So both the DSHB factorization (in Section 4) and the SDS factorization (in Section 5) are helpful in decentralized optimization, while the RHB factorization (in Section 3) is not suitable in this scenario.
@@ 7 Conclusion
In this paper, we study the sparse factorization $J=J_{0}AJ_{0}$, where $J=\tfrac{1}{n}\mathds{1}_{n}\mathds{1}_{n}^{\mathsf{T}}$ is the (scaled) all-ones matrix and $J_{0}=J_{1}\oplus\cdots\oplus J_{\tau}$ is the direct sum of several smaller all-ones matrices. We introduce the hierarchically banded structure of a symmetric matrix, based on which we present two types of hierarchically banded factorization of $J$: the reduced hierarchically banded (RHB) factorization and the doubly stochastic hierarchically banded (DSHB) factorization. Moreover, inspired by the DSHB factorization, we propose the sequential doubly stochastic (SDS) factorization which further factorizes the matrix $A$ as the product of a sequence of symmetric, doubly stochastic matrices. We then discuss the usefulness of the proposed factorizations in the decentralized average consensus problem and decentralized optimization. The presented three types of sparse factorization offer much flexibility in handling the trade-off between the per-iteration communication cost and the total number of communication rounds in decentralized averaging (and optimization).
Finally, recall that the partition $n=\sum_{k=1}^{\tau}n_{k}$ is assumed to be given and fixed throughout the paper. Further investigation is needed in the design of this partition to fully leverage the power of the proposed sparse factorizations in decentralized optimization.

### v2

@@ 6.2 Decentralized optimization
$\text{minimize}\ \ f(x):=\frac{1}{n}\sum\limits_{i=1}^{n}f_{i}(x),$ (29)
@@
where the optimization variable is $x\in\mathbb{R}^{d}$, and each component function $f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}$ is continuously differentiable and potentially nonconvex. Each agent $i\in\mathcal{V}$ only has access to one component function $f_{i}$, and agents communicate with each other via (time-varying) topologies $\{\mathcal{G}^{(k)}\}$. It can be shown that the decentralized average consensus problem is a special case of (29) with $f_{i}(x)=\tfrac{1}{2}\|x-x_{i}^{(0)}\|_{2}^{2}$.
@@
In the context of decentralized optimization, a sparse factorization of $J$ offers sequences of graphs that satisfy the finite-time consensus property, and incorporating such graph sequences in decentralized optimization algorithms could significantly reduce the communication cost in the algorithm while achieving a comparable convergence rate (compared with decentralized algorithms using traditional communication protocols) [21, 12]. Consider, for example, the decentralized gradient descent (DGD) algorithm, an extension of the gradient descent method to the decentralized setting:
$X^{(k+1)}=W^{(k)}X^{(k)}-\alpha\nabla\boldsymbol{f}(X^{(k)}),$
where $\nabla\boldsymbol{f}(X):=\begin{bmatrix}\nabla f_{1}(x_{1})&\cdots&\nabla f_{n}(x_{n})\end{bmatrix}^{\mathsf{T}}\in\mathbb{R}^{n\times d}$ is a matrix of all the component gradient functions and $\alpha\in\mathbb{R}_{>0}$ is the step size. At each DGD iteration, the matrix $W^{(k)}$ is the weighted adjacency matrix of a graph $\mathcal{G}^{(k)}$, and it can be sparse factors of $J$. More specifically, in DGD, one iterates from $W^{(1)}$ to $W^{(q)}$ in (2) and restart with $W^{(1)}$ again. In most existing analyses for DGD (and other decentralized optimization algorithms), the weight matrices $\{W^{(k)}\}$ are assumed to be connected and doubly stochastic (see, e.g., the book [13, §11.3]). Recall that the $T$-factors (21) are doubly stochastic yet not connected. So when the $T$-factors are used in DGD, the convergence guarantee is different from classical results. To see this, recall that the convergence rate of DGD with a connected static graph is inversely proportional to the so-called spectral gap $(1-\rho(W))$ [23], where $\rho(W)$ is the second largest eigenvalue (in modulus) of $W$. For time-varying connected graphs, the convergence rate is inversely proportional to the worst-case spectral gap $(1-\rho_{\mathrm{max}})$ [8], where $\rho_{\mathrm{max}}:=\max\{\rho(W^{(k)})\colon k\in\mathbb{N}\}$. In contrast, when $T$-factors are used, the convergence rate of DGD follows from [21, Theorem 1] (because the topology sequence constructed using (2) and $T$-factors satisfies all the assumptions stated in [21]) and reads as
$\frac{1}{K}\sum_{k=1}^{K}\|\nabla\boldsymbol{f}(X^{(k)})\|^{2}=O\left(\frac{nq^{2}}{K}\right),$ (30)
where $q$ is the finite-time consensus parameter in (1). Unlike the classical convergence results that depend on the spectral gap, the rate (30) is independent of the connectivity of any of the individual graphs, but instead considers the joint effect of all the topologies used in the algorithm. Therefore, the proposed factorization enables DGD to handle potentially sparser topologies that can be disconnected during certain iterations. The same conclusion can also be drawn for DGD with momentum [21] and the (decentralized) gradient tracking algorithm [12].
Numerical experiments
Numerical evidence is presented to demonstrate the potential benefits of using the $T$-factors in decentralized optimization algorithms. We apply DGD to solve the least squares problem (i.e., (29) with each $f_{i}(x)=\|A_{i}x-b_{i}\|^{2}$). The entries in each $A_{i}\in\mathbb{R}^{m\times d}$ are independently and identically distributed (IID) random variables drawn from the standard distribution, and so are the vectors $\{\tilde{x}_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$. The vector $b_{i}\in\mathbb{R}^{d}$ is then computed by $b_{i}=A_{i}\tilde{x}_{i}+\delta z_{i}$, where $\delta\in\mathbb{R}_{>0}$ is a prescribed constant and $z_{i}\in\mathbb{R}^{d}$ is Gaussian random noise. In the experiments, we set $n=241$ (a prime number), $m=100$, $d=50$, and $\delta=0.1$. We construct the partition (3) via binary representation of $n$, use the one-peer hyper-cubes in Phases 1 and 3, and use the proposed $T$-factors (21) in Phase 2.
Figure 3 presents the simulation results for various sparse topologies that have a maximum degree of $1$. We see that DGD with $T$-factors has similar, if not better, performance compared to DGD with other sparse topologies. Yet, only the proposed SDS factorization considers the clustering and hierarchical structures in modern application scenarios. So, despite the similar convergence rate, using the proposed $T$-factors takes fewer inter-cluster communication and has lower total communication cost. An intriguing observation is that the proposed SDS factorization with $T$-factors exhibits oscillation in the mean-square error (MSE) over time, whereas DGD with one-peer exponential graphs approaches a more stable MSE asymptotically. A closer examination finds that the oscillation period equals to $q$, the length of the factorization in (2). This suggests that the oscillation may stem from the three-phase nature of SDS factorization and the presence of some disconnected factors in SDS factorization. Consequently, during each period, agents across different clusters may have different iterates $x_{i}^{(k)}$ and only achieve consensus at the end of each period.
Figure 3: Mean-square error (MSE) versus the number of DGD iterations for the decentralized least squares problem, with various sparse graphs. The convergence of DGD with one-peer exponential graphs is plotted in blue, and that with one-peer undirected EquiTopo graphs [15] is in green.
@@ 7 Conclusion
In this paper, we study the sparse factorization $J=J_{0}AJ_{0}$, where $J=\tfrac{1}{n}\mathds{1}_{n}\mathds{1}_{n}^{\mathsf{T}}$ is the (scaled) all-ones matrix and $J_{0}=J_{1}\oplus\cdots\oplus J_{\tau}$ is the direct sum of several smaller (scaled) all-ones matrices. We introduce the hierarchically banded structure of a symmetric matrix, based on which we present two types of hierarchically banded factorization of $J$: the reduced hierarchically banded (RHB) factorization and the doubly stochastic hierarchically banded (DSHB) factorization. Moreover, inspired by the DSHB factorization, we propose the sequential doubly stochastic (SDS) factorization, which further factorizes the matrix $A$ as the product of a sequence of symmetric, doubly stochastic matrices. We then discuss the usefulness of the proposed factorizations in decentralized average consensus and decentralized optimization. The presented three types of sparse factorization offer much flexibility in handling the trade-off between the per-iteration communication cost and the total number of communication rounds in decentralized averaging (and optimization).
Finally, recall that the partition (3) is assumed to be given and fixed throughout the paper. Further investigation is needed in the design of this partition to fully leverage the power of the proposed sparse factorizations in decentralized optimization.